# Lifelong Knowledge Editing for LLMs with Retrieval-Augmented Continuous Prompt Learning

### v2

@@ 5.2 Efficiency Comparison
To underscore the efficiency of RECIPE, we conduct a comparative analysis on editing and inference time after 10,000 edits, as delineated in Table 3. Among methods leveraging edit-specific training such as MEND Mitchell et al. (2022), MALMEN Tan et al. (2023), LTE Jiang et al. (2024), and RECIPE, a notable reduction in editing time is observed when compared to techniques necessitating multiple iterations of back-propagation during editing. For inference speed, methods that modify model parameters maintain consistent speeds as they do not alter the original inference pipeline. T-Patcher Huang et al. (2023) slows down the inference speed due to the accumulating neurons. Among retrieval-based methods, GRACE Hartvigsen et al. (2022) reduces the parallelism in model inference due to its unique dictionary pairing mechanism. R-ROME Han et al. (2023) and LTE Jiang et al. (2024) need to calculate editing matrices on the fly and concatenate long editing instructions, respectively. In contrast, RECIPE effectively preserves the LLM’s original inference speed by concatenating short continuous prompts for editing. The shortest total time also highlights RECIPE’s efficiency advantage.
@@ 6 Conclusion
7 Acknowledgments
This work has been supported by Alibaba Group through Alibaba Innovative Research Program and Alibaba Research Intern Program.
@@ Appendix A Algorithms of RECIPE
Algorithm 1 Training of RECIPE
1:  Input: LLM to be edited $f_{llm}$; initialized collection of RECIPE parameters $\mathcal{M}$; training set $\mathcal{D}=\left\{\left((q_{e_{i}},a_{e_{i}}),\{q_{g_{i}}^{j},a_{g_{i}}^{j}\}_{j=1}^{N_{g_{i}}},\{q_{l_{i}}^{j},a_{l_{i}}^{j}\}_{j=1}^{N_{l_{i}}}\right)\right\}_{i=1}^{N}$; maximum iteration number $I_{max}$; batch size of training samples $b$; learning rate $\eta$.
2:  Output: trained RECIPE parameters $\mathcal{M}$.
3:  while $iter<I_{max}$ do
4:     $\left\{(q_{e_{i}},a_{e_{i}}),(q_{g_{i}},a_{g_{i}}),(q_{l_{i}},a_{l_{i}})\right\}_{i=1}^{b}\leftarrow$ Sample $b$ training samples from $\mathcal{D}$
5:     for $i\leftarrow 1$ to $b$ do
6:        # Get editing knowledge
7:        $k_{i}=q_{e_{i}}+a_{e_{i}}$ String concatenation
8:        # Get knowledge representation
9:        $r_{k_{i}}\leftarrow$ Transform $k_{i}$ using Eq. 6
10:        # Get continuous prompts
11:        $p_{k_{i}}\leftarrow$ Transform $r_{k_{i}}$ using Eq. 7
12:        # Get query representations
13:        $\tilde{r}_{q_{e_{i}}}\leftarrow$ Transform $q_{e_{i}}$ using Eq. 8
14:        $\tilde{r}_{q_{g_{i}}}\leftarrow$ Transform $q_{g_{i}}$ using Eq. 8
15:        $\tilde{r}_{q_{l_{i}}}\leftarrow$ Transform $q_{l_{i}}$ using Eq. 8
16:     end for
17:     # Get knowledge representation of KS
18:     $r_{\Theta}\leftarrow$ Transform $\Theta$ using Eq. 6
19:     # Compute loss and update parameters
20:     $\mathcal{L}_{edit},\mathcal{L}_{pl}\leftarrow$ Compute losses using Eq.13 and Eq.16
21:     $\mathcal{L}_{total}=\mathcal{L}_{edit}+\mathcal{L}_{pl}$
22:     $\mathcal{M}\leftarrow\operatorname{Adam}\left(\nabla_{\mathcal{M}}\mathcal{L}_{total},\eta\right)$
23:  end while
24:  return  $\mathcal{M}$
@@ B.3 Baselines
# Editing Editor ZSRE CF RIPE
@@
Table 5: The overall results using GPT-J (6B) in lifelong editing scenario.
Appendix C Model Settings and Training Details
RECIPE. (1) Hyper-parameter Settings: For our proposed RECIPE, we use the same hyper-parameter settings across different backbones, including LLAMA-2333https://huggingface.co/meta-llama/Llama-2-7b-hf, GPT-J [^4] https://huggingface.co/EleutherAI/gpt-j-6b, and GPT2-XL [^5] https://huggingface.co/openai-community/gpt2-xl. The number of continuous prompt tokens and KS tokens are set as $l=3$ and $c=10$, respectively. $\mathbf{MLP}_{K}$, $\mathbf{MLP}_{Q}$, and $\mathbf{MLP}_{P}$ are each composed of two linear layers, with an intermediate dimension set to 4096 and are connected in a residual manner He et al. (2016). The dimensions of the knowledge and query representations are also set to 4096. (2) Training Details: We set the learning rate ($\eta=1e-5$), the batch size to 8, and the maximum number of iterations to 150,000. A checkpoint is saved every 5000 iterations, and ultimately, the one with the smallest loss is selected for evaluation. The training process requires approximately 3 days on an NVIDIA A800 GPU. These experiments are presented on average with $5$ random runs, using different random seeds but the same hyper-parameters.
Appendix D Results with Different Backbones
Lifelong editing experiments on GPT-J (6B) and GPT2-XL (1.5B) are are presented in Table 5 and Table 6. The results also show a similar conclusion with general results, demonstrating the efficacy of our method.
# Editing Editor ZSRE CF RIPE
@@ Appendix D Results with Different Backbones
Table 6: The overall results using GPT2-XL (1.5B) in lifelong editing scenario.

### v3

@@ 5.2 Efficiency Comparison
To underscore the efficiency of RECIPE, we conduct a comparative analysis on editing and inference time after 10,000 edits, as delineated in Table 3. Among methods leveraging edit-specific training such as MEND, MALMEN, LTE, and RECIPE, a notable reduction in editing time is observed when compared to techniques necessitating multiple iterations of back-propagation during editing. For inference speed, methods that modify model parameters maintain consistent speeds as they do not alter the original inference pipeline. T-Patcher slows down the inference speed due to the accumulating neurons. Among retrieval-based methods, GRACE reduces the parallelism in model inference due to its unique dictionary pairing mechanism. R-ROME and LTE need to calculate editing matrices on the fly and concatenate long editing instructions, respectively. In contrast, RECIPE effectively preserves the LLM’s original inference speed by concatenating short continuous prompts for editing. The shortest total time also highlights RECIPE’s efficiency advantage.
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Key Research and Development Program of China under Grant No. 2022ZD0120302.
@@ B.2 General Datasets of LLMs
Among them, we use selection accuracy as the evaluation metric for CSQA, MMLU, and ANLI, and the reciprocal of the model’s perplexity (PPL) on the predicted sequence as the evaluation metric for SQuAD-2. All metrics are scaled to a range of 0-100.
@@ B.3 Baselines
Appendix C Model Settings and Training Details
RECIPE. (1) Hyper-parameter Settings: For our proposed RECIPE, we use the same hyper-parameter settings across different backbones, including LLAMA-$2$https://huggingface.co/meta-llama/Llama-2-7b-hf, GPT-J [^3] https://huggingface.co/EleutherAI/gpt-j-6b, and GPT2-XL [^4] https://huggingface.co/openai-community/gpt2-xl. The number of continuous prompt tokens and KS tokens are set as $l=3$ and $c=10$, respectively. $\mathbf{MLP}_{K}$, $\mathbf{MLP}_{Q}$, and $\mathbf{MLP}_{P}$ are each composed of two linear layers, with an intermediate dimension set to 4096 and are connected in a residual manner He et al. (2016). The dimensions of the knowledge and query representations are also set to 4096. The total numbers of RECIPE’s training parameters for GPT2-XL, GPT-J, and LLAMA-2 are 220M, 250M, and 250M, respectively. (2) Training Details: We set the learning rate ($\eta=1e-5$), the batch size to 8, and the maximum number of iterations to 150,000. A checkpoint is saved every 5000 iterations, and ultimately, the one with the smallest loss is selected for evaluation. The training process requires approximately 3 days on an NVIDIA A800 GPU. These experiments are presented on average with $5$ random runs, using different random seeds but the same hyper-parameters.
Appendix D Results with Different Backbones
Lifelong editing experiments on GPT-J (6B) and GPT2-XL (1.5B) are are presented in Table 5 and Table 6. The results also show a similar conclusion with general results, demonstrating the efficacy of our method. Notably, comparing Tables 1, 5, and 6, ROME and MEMIT exhibit a significant performance decline on LLAMA-2 compared to the other two backbones. MEMIT is an improved version of ROME, and both methods aim to edit the weights of the selected FFN layers. We speculate that this performance discrepancy may stem from structural differences in the FFN layers between LLAMA-2, GPT2-XL, and GPT-J. The original ROME and MEMIT papers assume a common two-layer structure for the LLM’s FFN layers, which is exactly the configuration of GPT2-XL and GPT-J. However, the FFN of LLAMA-2 consists of three linear layers, which may account for the ineffectiveness of ROME and MEMIT in adapting to the LLAMA-2 architecture.
# Editing Type Editor ZSRE CF RIPE
@@ Appendix D Results with Different Backbones
Table 5: The overall results using GPT-J (6B) in lifelong editing scenario. “# Editing” denotes the number of edits. “Rel.”, “Gen.” and “Loc.” are the abbreviations of reliability, generality, and locality, respectively. Given that the RIPE dataset comprises 4,388 samples, achieving results for 10,000 edits is not feasible. MP, AP, and RB indicate Modifying Parameters, Adding Parameters, and Retrieval-Based methods, respectively. The t-tests demonstrate the improvements of our work are statistically significant with $p$ < 0.05 level.
# Editing Type Editor ZSRE CF RIPE
@@
Table 6: The overall results using GPT2-XL (1.5B) in lifelong editing scenario. “# Editing” denotes the number of edits. “Rel.”, “Gen.” and “Loc.” are the abbreviations of reliability, generality, and locality, respectively. Given that the RIPE dataset comprises 4,388 samples, achieving results for 10,000 edits is not feasible. MP, AP, and RB indicate Modifying Parameters, Adding Parameters, and Retrieval-Based methods, respectively. The t-tests demonstrate the improvements of our work are statistically significant with $p$ < 0.05 level.
Appendix E Summary of Innovations
We summarize the innovations of this work as follows.
First, to the best of our knowledge, RECIPE is the first method to employ prompt learning for solving model editing. Indeed, previous works have used string-constructed prefixes to accomplish model editing, such as IKE Zheng et al. (2023) and LTE Jiang et al. (2024). However, we are the first to explore how to achieve model editing with the shortest editing prefixes through prompt learning. Second, a direct application of prompt learning might involve training a continuous prompt as a fixed prefix, which can append any editing string behind it to form a complete instruction for the LLM to follow. This approach treats the entire editing task as a prompt learning task. However, the concatenation of the continuous prompt with the editing string results in a lengthy prefix. Moreover, it can be reasonably hypothesized that a smaller amount of trainable parameters might make it challenging to converge to satisfactory editing performance. In contrast, our RECIPE innovatively treats the editing of each piece of knowledge as a mini prompt learning task, and considers the training of the encoder that generates continuous editing prompts as a meta-task. This strategy significantly reduces the length of editing prefixes while enhancing the representational capabilities of the editing training, thus ensuring editing performance. Third, RECIPE completely decouples knowledge editing from model parameters through prompt learning, thereby avoiding model degradation due to external intervention in model parameters or overfitting on the editing dataset.